# EFlat-LoRA: Efficiently Seeking Flat Minima for Better Generalization in Fine-Tuning Large Language Models and Beyond

### v1

@@ Abstract
Little research explores the correlation between the expressive ability and generalization ability of the low-rank adaptation (LoRA). Sharpness-Aware Minimization (SAM) improves model generalization for both Convolutional Neural Networks (CNNs) and Transformers by encouraging convergence to locally flat minima. However, the connection between sharpness and generalization has not been fully explored for LoRA due to the lack of tools to either empirically seek flat minima or develop theoretical methods. In this work, we propose Flat-LoRA and its efficient version i.e., EFlat-LoRA, to seek flat minima for LoRA. Concretely, we theoretically demonstrate that perturbations in the full parameter space can be transferred to the low-rank subspace. This approach eliminates the potential interference introduced by perturbations across multiple matrices in the low-rank subspace. Our extensive experiments on large language models and vision-language models demonstrate that EFlat-LoRA achieves optimize efficiency comparable to that of LoRA while simultaneously attaining comparable or even better performance. For example, on the GLUE dataset with RoBERTa-large, EFlat-LoRA outperforms LoRA and full fine-tuning by 1.0% and 0.5% on average, respectively. On vision-language models e.g., Qwen-VL-Chat shows performance improvements of 1.5% and 1.0% on SQA and VizWiz datasets, respectively. These empirical results also verify that the generalization of LoRA is closely related to sharpness, which is omitted by previous methods.
@@ Introduction
Parameter-Efficient Fine-Tuning (PEFT) methods only update a small subset of parameters, e.g., adapters (Houlsby et al. 2019)(Hu et al. 2022) or prompt weights (Li and Liang 2021)(Lester, Al-Rfou, and Constant 2021) for Large language models (LLMs) with substantially lower memory and computational costs. Specifically, Low-Rank Adaptation (LoRA) (Hu et al. 2022) stands out for achieving performance comparable to full fine-tuning (FT) while being considerably more efficient.
Figure 1: Comparison of Methods: LoRA, Flat-LoRA, and EFlat-LoRA.
@@
Many works have been proposed to enhance the performance of LoRA by introducing more dedicated budgets for rank allocation (Zhang et al. 2023b), decomposing optimization for direction and magnitude updates (Liu et al. 2024), or designing a better initialization strategy for LoRA parameters (Meng, Wang, and Zhang 2024)(Wang et al. 2024), etc. These studies demonstrate the significant potential for improving LoRA performance. However, most existing approaches fail to effectively address bias inheritance, where LLMs may propagate and amplify their inherent biases that significantly impact model performance and robustness on downstream tasks (Li et al. 2025). Therefore, a natural question is: how to model and understand the generalization of LoRA for various LLMs and beyond, e.g., vision-language models?
@@
It is widely believed that a flatter loss landscape can lead to better generalization performance (Hochreiter and Schmidhuber 1994) (Hochreiter and Schmidhuber 1997). For instance, Foret et al. proposed Sharpness-Aware Minimization (SAM) (Foret et al. 2021), which seeks parameter regions where the training loss remains uniformly flat. SAM and its variants have demonstrated State-Of-The-Art (SOTA) performances across various applications, such as classification (Zhou et al. 2023) (Kwon et al. 2021), transfer learning (Du et al. 2022) (Zhuang et al. 2022), domain generalization (Dong et al. 2024) (Xie, Li, and Wu 2024) and federated learning (Dai et al. 2023).
To our best knowledge, compared with theoretical analysis, e.g.(Neyshabur et al. 2017), empirically connecting between sharpness and generalization ability of LoRA is a practical approach, e.g.(Andriushchenko et al. 2023). For the second line of research, a naive method is to combine SAM with LoRA.
@@
However, if perturbations in SAM are applied simultaneously to two low-rank subspaces of LoRA, they may change the maximum loss within the neighborhood of LoRA’s full parameter space (Dinh et al. 2017a); besides, SAM incurs a computational cost twice that of Stochastic Gradient Descent (SGD) (Deng et al. 2024). Obviously, the core of question for the second line is how to efficiently seek flat minima of LoRA for empirically understanding of the connection between sharpness and generalization in LoRA.
In this paper, we propose a novel PEFT method, Flat-LoRA, that promotes convergence toward flatter minima. Specifically, we theoretically uncover that perturbations in the full parameter space can be equivalently re-parameterized as perturbations within the low-rank space.
In addition, we propose EFlat-LoRA to accelerate Flat-LoRA by an Exponential Moving Average (EMA) strategy. We validate that EFlat-LoRA improves the generalization performance on downstream tasks while maintaining computational efficiency comparable to that of LoRA. Fig.1 compares three methods: LoRA, Flat-LoRA, and EFlat-LoRA.
@@
We conducted comprehensive experiments on diverse tasks (fine-tuning, few-shot learning) and various model types (RoBERTa (Liu et al. 2019), GPT-2 (Radford et al. 2019), CLIP (Zanella and Ben Ayed 2024), Qwen-VL-Chat (Bai et al. 2023)) and scales. We find that our EFlat-LoRA can achieve model accuracy very close to or even surpass both full fine-tuning and LoRA across many tasks. Our main contribution can be summarized as follows:
@@
  • We propose Flat-LoRA, a novel PEFT training method that integrates SAM into the LoRA framework. Furthermore, the EFlat-LoRA provides an efficient tool for empirically understanding the connection between sharpness and generalization in LLMs and beyond in the future. We empirically show that reducing sharpness is highly correlated with improved generalization in PEFT tasks, which has been rarely explored in PEFT studies before.
@@
  • We conduct comprehensive experiments on LLMs (e.g., RoBERTa, GPT-2) and vision-language models (e.g., CLIP, Qwen-VL-Chat) across various tasks including fine-tuning and few-shot learning. Results show that EFlat-LoRA achieves optimize efficiency comparable to that of LoRA while simultaneously attaining comparable or even better performance.
@@
Hu et al. proposed LoRA (Hu et al. 2022) as a PEFT method that introduced low-rank adapters into each layer of a pre-trained model. Recent advancements in LoRA can be broadly categorized into two directions: 1) advanced architecture and 2) optimization method. In the first research line, for example, LoraHub (Huang et al. 2023) trained multiple adapters and strategically combined them based on the domain during inference. LoRA-FA (Zhang et al. 2023a) chose to freeze the projection-down weight of $\mathbf{A}$ and update the projection-up weight of $\mathbf{B}$ in each LoRA layer. DoRA (Liu et al. 2024) improved LoRA by incorporating a learnable magnitude vector to re-scale the normalized product of low-rank matrices. HydraLoRA (Tian et al. 2024) extended the LoRA framework with an asymmetric architecture that shared a common $\mathbf{A}$ matrix for efficiency while dynamically assigning samples to multiple $\mathbf{B}$ matrices via a MoE mechanism. In the second line, for example, LoRA+ (Hayou, Ghosh, and Yu 2024) applied different learning rates to the two low-rank matrices. Additionally, Galore (Zhao et al. 2024) employed SVD to compress the gradients and its first and second momentum of full training into a low-rank space, thereby reducing the memory footprint during pre-training and fine-tuning. Recently, Li et al. (Li et al. 2024a) proposed combining SAM with LoRA for better generalization, but they used random perturbation. Our method belongs to the second research line. Different from (Li et al. 2024a), our method transfers the perturbation from the full parameter space to a single low-rank parameter space without changing the maximum perturbed loss, avoiding misalignment with SAM’s training behavior.
@@
Research on the relationship between sharpness and generalization could be traced back to (Hochreiter and Schmidhuber 1997). Following the observation by (Keskar et al. 2017) that larger batch sizes tended to increase sharpness and generalization error, this relationship received substantial attention. (Jastrzkebski et al. 2017) extended this by finding a correlation between the sharpness and the ratio of learning rate to batch size. (Dinh et al. 2017b) showed that one can easily construct networks with good generalization but with arbitrary large sharpness by reparameterization. (Jiang et al. 2020) performed a large-scale empirical study on various generalization measures and showed that sharpness-based measures have the highest correlation with generalization. Theoretical understandings on the generalization error using sharpness-related measures were provided in (Dziugaite and Roy 2017), (Neyshabur et al. 2017), (Wang and Mao 2022). Collectively, these studies justified the goal of seeking flatter minima to improve generalization. However, to the best of our knowledge, the correlation between sharpness and generalization for LoRA has barely been discussed due to the lack of theoretical understanding or efficient tools for empirical analysis. Our method provides an efficient tool for empirical analysis in this area.
@@
Foret et al. (Foret et al. 2021) proposed the SAM to enhance model generalization as follows:
$\mathop{\min}\limits_{\mathbf{w}}[(\mathop{\max}\limits_{||\bm{\varepsilon}||\leq\rho}L(\mathbf{w}+\bm{\varepsilon})-L(\mathbf{w}))+L(\mathbf{w})+\lambda||\mathbf{w}||_{2}^{2}],$ (1)
where $\mathbf{w}$ represents the weights of the network, $\bm{\varepsilon}$ represents the perturbation of weights $\mathbf{w}$ in a Euclidean ball with the radius $\rho$ $(\rho>0)$, $L(\cdot)$ is the loss function, and $\lambda||\mathbf{w}||_{2}^{2}$ is a standard L2 regularization term.
SAM utilizes Taylor expansion to search for the maximum perturbed loss ($\mathop{\max}\limits_{||\bm{\varepsilon}||\leq\rho}L(\mathbf{w}+\bm{\varepsilon})$) in local parameter space as follows:
$\mathop{\arg\max}\limits_{||\bm{\varepsilon}||\leq\rho}\;L(\mathbf{w}+\bm{\varepsilon})\approx\mathop{\arg\max}\limits_{||\bm{\varepsilon}||\leq\rho}\;{\bm{\varepsilon}^{\top}}{\nabla_{\mathbf{w}}}L(\mathbf{w}).$ (2)
@@
$\hat{\bm{\varepsilon}}=\rho{\nabla_{\mathbf{w}}}L(\mathbf{w})/||{\nabla_{\mathbf{w}}}L(\mathbf{w})||.$ (3)
@@
${\nabla_{\mathbf{w}}}\mathop{\max}\limits_{||\bm{\varepsilon}||\leq\rho}L(\mathbf{w}+\bm{\varepsilon})\approx{\nabla_{\mathbf{w}}}L({\mathbf{w}}+\hat{\bm{\varepsilon}}({\mathbf{w}}))$ (4)
@@
${\nabla_{\mathbf{w}}}\mathop{\max}\limits_{||\bm{\varepsilon}||\leq\rho}L(\mathbf{w}+\bm{\varepsilon})\approx{\nabla_{\mathbf{w}}}L(\mathbf{w}){|_{\mathbf{w}+\bm{\hat{\varepsilon}}}}.$ (5)
@@
Recently, SAM variants could be broadly categorized into three groups: 1) studies on the perturbation radius $\rho$ in SAM, 2) studies that speed up the optimization process of SAM, and 3) redefinitions of sharpness in SAM. For the first direction, Kwon et al. (Kwon et al. 2021) proposed Adaptive SAM (ASAM), which adapted the perturbation radius in a scale-aware manner, allowing SAM to be effectively applied to scale-invariant neural networks. For the second group, Kim et al. (Kim et al. 2023) introduced a multi-step ascent approach to improve SAM. Li et al. (Li et al. 2024b) introduced Friendly SAM (F-SAM), which improved generalization by removing the detrimental influence of the full gradient component and instead utilizing batch-specific gradients to guide optimization more effectively. For the third group, Zhuang et al. (Zhuang et al. 2022) pointed out that SAM did not always favor flat minima. Consequently, they proposed GSAM, which minimized the surrogate gap and the perturbed loss to better encourage flatness. Zhang et al. introduced the first-order flatness (Zhang et al. 2023c), which assessed the maximal gradient norm within a perturbation radius. Consequently, they proposed GAM which explicitly seeks minima characterized by uniformly small curvature.
@@
LoRA achieves parameter efficiency by modeling the low-rank decomposed weight (Li et al. 2018) (Li et al. 2022). Specifically, the weight change for each layer $\mathbf{W}_{0}\in\mathbb{R}^{n\times m}$ is represented as $\Delta\mathbf{W}=s\mathbf{B}\mathbf{A}$, where $s$ is a scaling factor, $\mathbf{B}\in\mathbb{R}^{n\times r}$, $\mathbf{A}\in\mathbb{R}^{r\times m}$, with rank $r\ll\min(n,m)$. Given input $\mathbf{x}$, the forward pass is as follows:
@@
where matrix $\mathbf{A}$ is typically initialized by the Kaiming’s method (He et al. 2015), $\mathbf{B}$ is set to zeros. $\mathbf{W}_{0}$ remains unchanged during the fine-tuning process, and the parameters $\mathbf{B}$ and $\mathbf{A}$ are being trained. At inference time, $\Delta\mathbf{W}$ is merged into $\mathbf{W_{0}}$.
@@
$\min_{\mathbf{A},\mathbf{B}}~~\mathop{\max}\limits_{\scriptstyle||{{\bf{E}}^{\bf{A}}}|{|_{F}}\leq\rho,\hfill\atop\scriptstyle||{{\bf{E}}^{\bf{B}}}|{|_{F}}\leq\rho\hfill}L({\mathbf{W_{0}}}+{s}(\mathbf{B}+{\mathbf{E}^{\mathbf{B}}})(\mathbf{A}+{\mathbf{E}^{\mathbf{A}}})),$ (7)
@@
$\hat{\bm{\varepsilon}}^{\mathbf{w}}=\rho\text{sign}(\mathbf{g}^{\mathbf{w}})\frac{\mathbf{g}^{\mathbf{w}}}{||\mathbf{g}^{\mathbf{w}}||},$ (9)
@@
where $(\mathbf{A}^{\top})^{+}$ and $(\mathbf{B}^{\top})^{+}$ represent the pseudo-inverse of $\mathbf{A}^{\top}$ and $\mathbf{B}^{\top}$, respectively. To obtain a more accurate estimate of the gradient of the full weights, we combine the above two approaches to compute $\nabla L_{\mathbf{W}}(\mathbf{W})$ as follows:
@@
Let ${\bar{\mathbf{g}}^{\mathbf{W}}}=\text{Vector}(\overline{\nabla{L}_{\mathbf{W}}(\mathbf{W})})$. Then the perturbation in Eq. (9) could be rewritten as fallows:
$\bar{\mathbf{E}}^{\mathbf{W}}=\text{Matrix}(\rho\text{sign}({\bar{\mathbf{g}}^{\mathbf{W}}})\frac{{\bar{\mathbf{g}}^{\mathbf{W}}}}{||{\bar{\mathbf{g}}^{\mathbf{W}}}||}),$ (13)
@@
where $\mathbf{A}^{+}$ is the pseudo-inverse of $\mathbf{A}$. An alternative approach is to transfer the perturbation to matrix $\mathbf{A}$. Following the observations from HydraLoRA (Tian et al. 2024), matrix $\mathbf{A}$ shows high parameter similarity across heads, likely due to initialization, making it capture domain-common features, while matrix $\mathbf{B}$ remains distinct and domain-specific. Since different tasks require different perturbations, we adopt the approach of transferring the perturbation to the matrix $\mathbf{B}$, as expressed in Eq. (14). Section A.1 and Section C in the supplementary file supplies the detailed derivation of Eq. (10) and the pseudo algorithm of Flat-LoRA.
Balancedness of Flat-LoRA.
Balancedness is well-appreciated in domains such as matrix factorization/sensing (Ge, Jin, and Zheng 2017) (Bartlett, Helmbold, and Long 2018) (Du, Hu, and Lee 2018). It is also observed that balanced neural networks are easier to optimize relative to unbalanced ones (Neyshabur, Salakhutdinov, and Srebro 2015). Recently, Balancedness $B_{t}:=\frac{1}{2}(||\mathbf{x}_{t}||^{2}-||\mathbf{y}_{t}||^{2})$ (where $\mathbf{x}_{t}$ and $\mathbf{y}_{t}$ are variables) turns out to be an intriguing alternative to sharpness on the scale-invariant problem (Li, Zhang, and He 2024).
To investigate the balancedness of our proposed method, we express the update process of Flat-LoRA analogously to Eq.(4) in (Li, Zhang, and He 2024) as follows:
@@
Theorem 1
Let $B_{t}:=\frac{1}{2}(||\mathbf{x}_{t}||^{2}-||\mathbf{y}_{t}||^{2})$. For the learning rate $\eta\Rightarrow 0$, the limiting flow of Flat-LoRA guarantees that:
$\left|{\frac{1}{2}\frac{{d({{\left\|{{\mathbf{x}_{t}}}\right\|}^{2}}-{{\left\|{{\mathbf{y}_{t}}}\right\|}^{2}})}}{{dt}}}\right|\leq\left|{\rho\frac{1}{s}\frac{1}{{\left\|{\mathbf{y}_{t}}\right\|}}\left\|{{\mathbf{g}_{{{{\rm{\tilde{\mathbf{x}}}}}_{t}}}}}\right\|}\right|.$ (17)
Theorem 1 indicates that the balancedness of Flat-LoRA is influenced by the perturbation range $\rho$, the norm of the gradient with $\mathbf{y}_{t}$ at the perturbed point, the $\ell_{2}$-norm of $\mathbf{y}_{t}$, and the scale constraint of LoRA. To ensure that the balancedness of Flat-LoRA gradually decreases during training, we progressively reduce $\rho$ throughout the training process. In addition, the norm of the gradient with respect to $\mathbf{y}_{t}$ at the perturbed point also decreases due to the weight decay. The $\ell_{2}$-norm of $\mathbf{y}_{t}$ is bounded within a certain range, these factors collectively contribute to the reduction in the balancedness of Flat-LoRA.
Figure 2: Parameter update process for EFlat-LoRA.
Efficient Flat-LoRA
The optimization processes of Flat-LoRA and SAM are identical, as both require two gradient computations within an iteration. To enhance optimization efficiency, we propose Efficient Flat-LoRA (EFlat-LoRA), which estimates the subsequent perturbation ${\mathbf{E}}^{\mathbf{B}}$ in Eq. (15) by maintaining an Exponential Moving Average (EMA) of previous perturbations as follows:
@@
where $\beta\in(0,1)$ is the momentum coefficient that determines the update rate of the exponential moving average. ${\mathbf{E}^{\mathbf{B}}_{t}}$ is the perturbation on matrix $\mathbf{B}_{t}$ at $t$-th iteration, ${{\hat{\mathbf{E}}}^{\mathbf{B}}_{t}}$ is the EMA perturbation at $t$-th iteration. Fig. 2 illustrates the parameter update process of EFlat-LoRA: (1) Calculate the gradient at the perturbed point $(\mathbf{W}_{0}$, $\mathbf{B}_{t-1}+\hat{\mathbf{E}}^{\mathbf{B}}_{t-1}$, $\mathbf{A}_{t-1})$. (2) Calculate the perturbation ${\mathbf{E}}^{\mathbf{B}}_{t}=\frac{1}{s}\bar{\mathbf{E}}^{\mathbf{W}}\mathbf{A}^{+}_{t-1}$. (3) Return to the original parameter point $(\mathbf{W}_{0},\mathbf{B}_{t-1},\mathbf{A}_{t-1})$. (4) Update the parameters to $(\mathbf{W}_{0},\mathbf{B}_{t},\mathbf{A}_{t})$. (5) Calculate the EMA perturbation by Eq.(18) and update the parameters to the next perturbed point $(\mathbf{W}_{0},\mathbf{B}_{t}+\hat{\mathbf{E}}^{\mathbf{B}}_{t},\mathbf{A}_{t})$. During this optimization process, each optimization step requires only a single forward and backward. The algorithmic pseudocode is provided in Section C of Supplementary file.
RoBERTa SST-2 SST-5 SNLI MNLI RTE TREC avg.$\uparrow$

### v2

# Efficiently Seeking Flat Minima for Better Generalization in Fine-Tuning Large Language Models and Beyond

@@ Abstract
Little research explores the correlation between the expressive ability and generalization ability of the low-rank adaptation (LoRA). Sharpness-Aware Minimization (SAM) improves model generalization for both Convolutional Neural Networks (CNNs) and Transformers by encouraging convergence to locally flat minima. However, the connection between sharpness and generalization has not been fully explored for LoRA due to the lack of tools to either empirically seek flat minima or develop theoretical methods. In this work, we propose Flat Minima LoRA (FMLoRA) and its efficient version i.e., EFMLoRA, to seek flat minima for LoRA. Concretely, we theoretically demonstrate that perturbations in the full parameter space can be transferred to the low-rank subspace. This approach eliminates the potential interference introduced by perturbations across multiple matrices in the low-rank subspace. Our extensive experiments on large language models and vision-language models demonstrate that EFMLoRA achieves optimize efficiency comparable to that of LoRA while simultaneously attaining comparable or even better performance. For example, on the GLUE dataset with RoBERTa-large, EFMLoRA outperforms LoRA and full fine-tuning by 1.0% and 0.5% on average, respectively. On vision-language models e.g., Qwen-VL-Chat, there are performance improvements of 1.5% and 1.0% on the SQA and VizWiz datasets, respectively. These empirical results also verify that the generalization of LoRA is closely related to sharpness, which is omitted by previous methods.
@@ Introduction
Parameter-Efficient Fine-Tuning (PEFT) methods only update a small subset of parameters, e.g., adapters (hu2022lora) or prompt weights (li2021prefix) for Large language models (LLMs) with substantially lower memory and computational costs. Specifically, Low-Rank Adaptation (LoRA) (hu2022lora) stands out for achieving performance comparable to full fine-tuning (FT) while being considerably more efficient.
Figure 1: Comparison of Methods: LoRA, FMLoRA, and EFMLoRA.
@@
Many works have been proposed to enhance the performance of LoRA by introducing more dedicated budgets for rank allocation (zhang2023adaptive), decomposing optimization for direction and magnitude updates (liu2024dora), or designing better initialization strategies for LoRA parameters (meng2024pissa), etc. These studies demonstrate the significant potential to improve LoRA performance. However, most existing approaches fail to effectively address bias inheritance, where LLMs may propagate and amplify their inherent biases, significantly impacting model performance and robustness on downstream tasks (li2025understanding). Therefore, a natural question is: how to model and understand the generalization of LoRA for various LLMs and beyond, e.g., vision-language models?
@@
It is widely believed that a flatter loss landscape can lead to better generalization performance (hochreiter1994simplifying) (hochreiter1997flat). For instance, Foret et al. proposed Sharpness-Aware Minimization (SAM) (foret-2020-SAM-ICLR), which seeks parameter regions where the training loss remains uniformly flat. SAM and its variants have demonstrated State-Of-The-Art (SOTA) performances across various applications, such as classification (kwon-2021-asam-ICML), transfer learning (zhuang-2022-GSAM-ICLR), domain generalization (dong2024implicit) and federated learning (FedGAMMA).
@@
To the best of our knowledge, compared to theoretical analysis, e.g.(neyshabur2017exploring), empirically connecting sharpness and generalization ability of LoRA is a practical approach, e.g.(andriushchenko2023modern). For the second line of research, a naive approach is to combine SAM with LoRA. However, if perturbations in SAM are applied simultaneously to two low-rank subspaces of LoRA, they may change the maximum loss within the neighborhood of LoRA’s full parameter space (dinh2017sharp); besides, SAM incurs a computational cost twice that of Stochastic Gradient Descent (SGD) (deng2024effective). The key question in the second line of research is how to efficiently find flat minima in LoRA, aiming to better understand the connection between sharpness and generalization.
@@
In this paper, we propose a novel PEFT method, FMLoRA, that promotes convergence toward flatter minima. Specifically, we theoretically uncover that perturbations in the full parameter space can be equivalently re-parameterized as perturbations within the low-rank space. In addition, we propose EFMLoRA to accelerate FMLoRA by an Exponential Moving Average (EMA) strategy. We validate that EFMLoRA improves generalization performance on downstream tasks while maintaining computational efficiency comparable to that of LoRA. Fig.1 compares three methods: LoRA, FMLoRA, and EFMLoRA. We conducted comprehensive experiments on diverse tasks (fine-tuning, few-shot learning) and various model types (RoBERTa (liu2019roberta), GPT-2 (radford2019language), CLIP (zanella2024low), Qwen-VL-Chat (Bai2023QwenVLAV)) and scales. We find that EFMLoRA achieves model accuracy very close to, or even surpass both full fine-tuning and LoRA across many tasks. Our main contribution can be summarized as follows:
@@
  • We propose FMLoRA, a novel PEFT training method that integrates SAM into the LoRA framework. Furthermore, EFMLoRA provides an efficient tool for empirically understanding the connection between sharpness and generalization in LLMs and beyond. We empirically show that reducing sharpness is highly correlated with improved generalization in PEFT tasks, which has been rarely explored in PEFT studies before.
@@
  • We conduct comprehensive experiments on LLMs (e.g., RoBERTa, GPT-2) and vision-language models (e.g., CLIP, Qwen-VL-Chat) across various tasks including fine-tuning and few-shot learning. Results show that EFMLoRA achieves optimize efficiency comparable to that of LoRA while simultaneously attaining comparable or even better performance.
@@
Hu et al. proposed LoRA (hu2022lora) as a PEFT method that introduced low-rank adapters into each layer of a pre-trained model. Recent advancements in LoRA can be broadly categorized into two directions: 1) advanced architectures and 2) optimization methods. In the first research line, for example, LoraHub (huang2023lorahub) trained multiple adapters and strategically combined them based on the domain during inference. LoRA-FA (zhang2023lora) chose to freeze the projection-down weight of $\mathbf{A}$ and update the projection-up weight of $\mathbf{B}$ in each LoRA layer. DoRA (liu2024dora) improved LoRA by incorporating a learnable magnitude vector to re-scale the normalized product of low-rank matrices. HydraLoRA (tian2024hydralora) extended the LoRA framework with an asymmetric architecture that shared a common $\mathbf{A}$ matrix for efficiency while dynamically assigning samples to multiple $\mathbf{B}$ matrices via a MoE mechanism. In the second line, for example, LoRA+ (hayou2024lora+) applied different learning rates to the two low-rank matrices. Additionally, Galore (zhao2024galore) employed SVD to compress the gradients and its first and second momentum of full training into a low-rank space, thereby reducing the memory footprint during pre-training and fine-tuning. Recently, Li et al. (li2024flat) proposed combining SAM with LoRA for better generalization, but they used random perturbation. Our method belongs to the second research line. Different from (li2024flat), our method transfers the perturbation from the full parameter space to a single low-rank parameter space without changing the maximum perturbed loss, avoiding misalignment with SAM’s training behavior.
@@
Research on the relationship between sharpness and generalization could be traced back to (hochreiter1997flat). Following the observation by (keskar-2016-large_batch-ICLR) that larger batch sizes tended to increase sharpness and generalization error. (jastrzkebski2017three) extended this by finding a correlation between the sharpness and the ratio of learning rate to batch size. (dinh-2017-sharp_minima-ICML) showed that one can easily construct networks with good generalization but with arbitrary large sharpness by reparameterization. (jiang-2019-fantastic-ICLR) performed a large-scale empirical study on various generalization measures and showed that sharpness-based measures have the highest correlation with generalization. Theoretical understandings on the generalization error using sharpness-related measures were provided in (neyshabur2017exploring), (wanggeneralization). Collectively, these studies justified the goal of seeking flatter minima to improve generalization. However, to the best of our knowledge, the correlation between sharpness and generalization for LoRA has barely been discussed due to the lack of theoretical understanding or efficient tools for empirical analysis. Our method provides an efficient tool for empirical analysis in this domain.
@@
Foret et al. (foret-2020-SAM-ICLR) proposed the SAM to enhance model generalization as follows:
[…]
@@
[…]
@@
[…]
@@
[…]
@@
Recently, SAM variants could be broadly categorized into three groups: 1) studies on the perturbation radius $\rho$ in SAM, 2) studies that speed up the optimization process of SAM, and 3) redefinitions of sharpness in SAM. For the first direction, Kwon et al. (kwon-2021-asam-ICML) proposed Adaptive SAM (ASAM), which adapted the perturbation radius in a scale-aware manner, allowing SAM to be effectively applied to scale-invariant neural networks. For the second group, Kim et al. (kim2023exploring) introduced a multi-step ascent approach to improve SAM. Li et al. (li2024friendly) introduced Friendly SAM (F-SAM), which improved generalization by removing the detrimental influence of the full gradient component and instead utilizing batch-specific gradients to guide optimization more effectively. For the third group, Zhuang et al. (zhuang-2022-GSAM-ICLR) pointed out that SAM did not always favor flat minima. Consequently, they proposed GSAM, which minimized the surrogate gap and the perturbed loss to better encourage flatness. Zhang et al. introduced the first-order flatness (zhang-2023-gradient-CVPR), which assessed the maximal gradient norm within a perturbation radius. Consequently, they proposed GAM which explicitly seeks minima characterized by uniformly small curvature.
@@
LoRA achieves parameter efficiency by modeling the low-rank decomposed weight (li2022low). Specifically, the weight change for each layer $\mathbf{W}_{0}\in\mathbb{R}^{n\times m}$ is represented as $\Delta\mathbf{W}=s\mathbf{B}\mathbf{A}$, where $s$ is a scaling factor, $\mathbf{B}\in\mathbb{R}^{n\times r}$, $\mathbf{A}\in\mathbb{R}^{r\times m}$, with rank $r\ll\min(n,m)$. Given an input $\mathbf{x}$, the forward is as follows:
@@
where matrix $\mathbf{A}$ is typically initialized by the Kaiming’s method (he2015delving), $\mathbf{B}$ is set to zeros. $\mathbf{W}_{0}$ remains unchanged during fine-tuning, while $\mathbf{B}$ and $\mathbf{A}$ are trained. During inference, $\Delta\mathbf{W}$ is merged into $\mathbf{W_{0}}$.
@@
[…]
@@
where $\mathbf{g}^{\mathbf{w}}=\text{Vector}(\nabla L_{\mathbf{W}}(\mathbf{W}))$ and $\hat{\bm{\varepsilon}}^{\mathbf{w}}=\text{Vector}(\hat{\mathbf{E}}^{\mathbf{W}})$, in which the $\text{Vector}(\cdot)$ function represents a vectorized operation. However, the solution for $\hat{\mathbf{E}}^{\mathbf{W}}$ explicitly depends on the gradient of the matrix $\mathbf{W}$. That is, the form of solution in Eq. (9) is undesirable since $\nabla L_{\mathbf{W}}(\mathbf{W})$ is unknown during LoRA optimization.
@@
where $(\mathbf{A}^{\top})^{+}$ and $(\mathbf{B}^{\top})^{+}$ represent the pseudo-inverse of $\mathbf{A}^{\top}$ and $\mathbf{B}^{\top}$, respectively. The accuracy of the pseudo-inverse depends on the condition number of matrix. A smaller condition number leads to a more accurate pseudo-inverse. Matrices with lower condition numbers are better suited for stable representation. In LoRA, we found that the condition number is typically low, around 3.
To obtain a more accurate estimate of the gradient of the full weights, we combine the above two approaches to compute $\nabla L_{\mathbf{W}}(\mathbf{W})$ as follows:
@@
Let ${\bar{\mathbf{g}}^{\mathbf{W}}}=\text{Vector}(\overline{\nabla{L}_{\mathbf{W}}(\mathbf{W})})$. Then the perturbation in Eq. (9) could be rewritten as follows:
@@
where $\text{Matrix}(\cdot)$ denotes the operation that converts a vector into a matrix. We transfer the perturbation from the full parameter space to a single low-rank parameter space without changing the maximum loss in the local region of the parameters. We apply no perturbation to matrix $\mathbf{A}$, i.e., ${\mathbf{E}^{\mathbf{A}}}=\mathbf{0}$, and ensure that the loss under perturbations in the low-rank subspace in Eq. (7) matches the inner maximum loss in Eq. (8), as follows:
@@
where $\mathbf{A}^{+}$ is the pseudo-inverse of $\mathbf{A}$. An alternative approach is to transfer the perturbation to matrix $\mathbf{A}$. Following the observations from HydraLoRA (tian2024hydralora), matrix $\mathbf{A}$ shows high parameter similarity across heads, likely due to initialization, making it capture domain-common features, while matrix $\mathbf{B}$ remains distinct and domain-specific. Since different tasks require different perturbations, we adopt the approach of transferring the perturbation to the matrix $\mathbf{B}$, as expressed in Eq. (14). The detailed derivation of Eq. (10) and the pseudo-algorithm for FMLoRA are provided in the supplementary file.
Balancedness of FMLoRA.
[…]
To investigate the balancedness of our proposed method, we express the update process of FMLoRA analogously to Eq.(4) in (li2024implicit) as follows:
@@
Theorem 1.
[…]
Theorem 1 indicates that the balancedness of FMLoRA is influenced by the perturbation range $\rho$, the norm of the gradient at the perturbed point, the $\ell_{2}$-norm of $\mathbf{y}_{t}$, and the scale constraint of LoRA. To ensure that the balancedness of FMLoRA gradually decreases during training, we reduce $\rho$ progressively. In addition, the norm of the gradient with respect to $\mathbf{y}_{t}$ at the perturbed point also decreases due to the weight decay. The $\ell_{2}$-norm of $\mathbf{y}_{t}$ is bounded within a certain range, these factors collectively contribute to the reduction in the balancedness of FMLoRA.
Figure 2: Parameter update process for EFMLoRA.
Efficient FMLoRA
The optimization processes of FMLoRA also require two gradient computations per iteration. To enhance optimization efficiency, we propose Efficient FMLoRA (EFMLoRA), which estimates the subsequent perturbation ${\mathbf{E}}^{\mathbf{B}}$ in Eq. (15) by maintaining an Exponential Moving Average (EMA) of previous perturbations as follows:
@@
where $\beta\in(0,1)$ is the momentum coefficient that determines the update rate of the exponential moving average. $\mathbf{E}^{\mathbf{B}}_{t}$} is the perturbation on matrix $\mathbf{B}_{t}$ at $t$-th iteration, ${\hat{\mathbf{E}}}^{\mathbf{B}}_{t}$} is the EMA perturbation at $t$-th iteration. Fig. 2 illustrates the parameter update process of EFMLoRA: (1) Calculate the gradient at the perturbed point $(\mathbf{W}_{0}$, $\mathbf{B}_{t-1}+\hat{\mathbf{E}}^{\mathbf{B}}_{t-1}$, $\mathbf{A}_{t-1})$. (2) Calculate the perturbation ${\mathbf{E}}^{\mathbf{B}}_{t}=\frac{1}{s}\bar{\mathbf{E}}^{\mathbf{W}}\mathbf{A}^{+}_{t-1}$. (3) Return to the original parameter point $(\mathbf{W}_{0},\mathbf{B}_{t-1},\mathbf{A}_{t-1})$. (4) Update the parameters to $(\mathbf{W}_{0},\mathbf{B}_{t},\mathbf{A}_{t})$. (5) Calculate the EMA perturbation by Eq.(18) and update the parameters to the next perturbed point $(\mathbf{W}_{0},\mathbf{B}_{t}+\hat{\mathbf{E}}^{\mathbf{B}}_{t},\mathbf{A}_{t})$. During this optimization process, each optimization step requires only a single forward and backward. The algorithmic pseudocode is provided in the supplementary file.
[…]